OpenStudy (anonymous):

anyone have a clue 9/2x-7/2=11/2

OpenStudy (anonymous):

9x/2 = 11/2 + 7/2 9x/2 = 9 Can you find x from it??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x = 18/8 Can now you find x or not??

OpenStudy (anonymous):

Solving an equation is like unwrapping a present. The present is the variable x. First, you untie all the ribbons and wrapping paper to get to the x. The ribbon is the -7/2. You get rid of it by adding 7/2 to both sides with yields 9/2*x=9 Now, the final wrapping paper is the 9/2 bit. You want to have x alone. So to get rid of the 9/2, you multiply both sides by 2/9 ('cuz 9/2 times 2/9 = 1 and that would get rid of the 9/2) and you get x=2. :D